Direct Materials and Direct Labor Variances

At the beginning of June, Bezco Toy Company budgeted 24,000 toy action figures to be manufactured in June at standard direct materials and direct labor costs as follows:

Direct materials $28,800
Direct labor 12,000
Total $40,800

The standard materials price is $0.4 per pound. The standard direct labor rate is $10 per hour. At the end of June, the actual direct materials and direct labor costs were as follows:

Actual direct materials $26,700
Actual direct labor 11,100
Total $37,800

There were no direct materials price or direct labor rate variances for June. In addition, assume no changes in the direct materials inventory balances in June. Bezco Toy Company actually produced 21,600 units during June.

Determine the direct materials quantity and direct labor time variances. Round your per unit computations to two decimal places, if required. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials quantity variance $ Unfavorable
Direct labor time variance $ Unfavorable
Answer #1

When there is no material price and rate variance then difference between standard cost and actual cost are for Material quantity and labor time variance

Direct material quantity variance = (28800/24000*21600)-26700 = 780 Unfavorable

Direct labor time variance = (12000/24000*21600)-11100 = 300 Unfavorable